# Exceptional Points in the Scattering Resonances of a Sphere Dimer

### v1

@@ Abstract
We investigate exceptional points of degeneracy (EPDs) in electromagnetic scattering of a sphere dimer from the electroquasistatic limit to the fully retarded regime. In the quasistatic limit, we prove that $\mathscr{P}\mathscr{T}$-symmetric configurations, realized by spheres with complex-conjugate susceptibilities, host EPDs. Beyond this limit, retardation breaks $\mathscr{P}\mathscr{T}$-symmetry; nevertheless, by jointly tuning the material dispersion of the two spheres, we derive analytic conditions for the existence of EPDs at real-frequencies. Near an EPD, we show that single-parameter perturbations yield the characteristic square-root splitting of the eigenfrequencies, and we quantify its impact on scattering, extinction, and absorption, clarifying sensing implications.
@@ I Introduction
Several studies [6, 7, 8] have employed coupled-mode theory (CMT) [13] to investigate EPDs in parity-time ($\mathscr{P}\mathscr{T}$) symmetric systems. While CMT is attractive because it is conceptually simple, offers clear physical intuition, and requires minimal computational effort, it does not precisely represent the underlying physics. In particular, coupled electromagnetic resonators are intrinsically non-$\mathscr{P}\mathscr{T}$ symmetric due to retardation in the radiative coupling, which leads to a complex and frequency-dependent coupling, that cannot be easily included in conventional CMT. Moreover, a recent study [14], employing the coupled-dipole approximation (CDA) [15, 16, 17, 18, 19, 20, 21], has shown that these retardation effects can themselves induce EPDs even in systems with no gain and loss.
In this work, we investigate EPDs in electromagnetic scattering by a dimer of homogeneous, linear, time-dispersive spheres with arbitrary material dispersion. We cast the problem into an equivalent two-port circuit, formally equivalent to the coupled-dipole approximation, but with the additional advantage of disentangling the dependencies on the excitation, geometry, and material properties, which is crucial for simplifying the analytical derivation and interpretation of EPD conditions. We prove that, in the electroquasistatic regime, a $\mathscr{P}\mathscr{T}$-symmetric sphere dimer with complex-conjugate susceptibilities supports EPDs. Beyond this limit, retardation breaks $\mathscr{P}\mathscr{T}$-symmetry; nevertheless, real-frequency EPDs can still emerge when the dispersion of the two materials are jointly tuned, and we derive general conditions for their occurrence. These predictions are validated against rigorous multiparticle Mie theory calculations [22]. Finally, we quantify the square-root spectral splitting to single-parameter perturbations and its impact on scattering, extinction, and absorption, showing that this platform enables sensors with great sensitivity.
@@ II The Model
Figure 1: (a) Dimer composed by two spheres of radius $a$ and inter-center distance $d$. (b) Two-port equivalent circuit of two interacting particles in the coupled dipole approximation and in the full-wave regime. (b) Two-port equivalent circuit of two interacting particles in the coupled dipole approximation and in the static limit. (c) Configuration for characterizing the resonant mode of the dimer (i.e., $Z_{\mathrm{in}}=0$). In panels (b),(c),(d) $h\in\{x,y,z\}$ denotes the polarization of the dipolar mode.
Consider a dimer composed of two spherical particles, labeled by $\mu=1,2$, each made by a linear, homogeneous, isotropic, and time-dispersive material, characterized by an electric susceptibility $\chi_{\mu}$, embedded in vacuum. Each sphere has radius $a$ and is centered at $\mathbf{r}_{\mu}$, with the inter-center distance denoted by $d$. A time-harmonic dependence $e^{j\omega t}$ is assumed. Let $\mathcal{B}_{\mu}$ be the bounded three-dimensional domain occupied by the $\mu$-th particle, and $\partial\mathcal{B}_{\mu}$ its boundary. Without loss of generality, we assume the dimer is aligned along the $x$-axis [see Fig. 1(a)].
The full-wave electromagnetic scattering problem can be formulated considering as unknown the polarization current density field $\mathbf{J}$ induced in the region $\mathcal{B}=\mathcal{B}_{1}\cup\mathcal{B}_{2}$ by the incident field $\mathbf{E}^{\mathrm{i}}$. In the following, we adopt the notation described in Appendix A, where for any vector field $\mathbf{F}=\mathbf{F}(\mathbf{r})$ we denote by $\mathbf{F}_{\mu}$ its restriction on $\mathcal{B}_{\mu}$. In the region $\mathcal{B}_{\mu}$, the field ${\bf J}_{\mu}$ is related to the total electric field ${\mathbf{E}}_{\mu}$, which is the sum of the scattered field ${\mathbf{E}}_{\mu}^{\mathrm{s}}$ and the incident field ${\mathbf{E}}_{\mu}^{\mathrm{i}}$, i.e. $\mathbf{E}_{\mu}=\mathbf{E}_{\mu}^{\mathrm{i}}+\mathbf{E}_{\mu}^{\mathrm{s}}$, by the relation
@@
where $Z_{\mu}^{\mathrm{m}}$ is the material impedance of the $\mu$-th particle [23], defined as
@@
where $\mathbf{E}^{\mathrm{s}}_{\nu,\mu}$ denotes the contribution to the scattered field on $\mathcal{B}_{\mu}$ arising from the current density $\mathbf{J}_{\nu}$ on the $\nu$-th particle alone. The scattered field $\mathbf{E}^{\mathrm{s}}_{\nu,\mu}$ is related to its source $\mathbf{J}_{\nu}$ through the integral operator $\mathscr{L}_{\mu}$ as
@@
where $I_{\mu}^{k}=I_{\mu}^{k}(\omega)$ are the weights. The current density modes $\mathbf{j}^{k}_{\mu}$ are the solutions of the following eigenvalue problem on $\mathcal{B}_{\mu}$:
$\mathscr{L}_{\mu}\left\{\mathbf{j}^{k}_{\mu}\right\}=\mathcal{Z}^{k}\,\mathbf{j}^{k}_{\mu},$ (6)
where $\mathcal{Z}^{k}=\mathcal{R}^{k}+j\mathcal{X}^{k}$ is the modal impedance of the current eigenmodes $\mathbf{j}_{\mu}^{k}$, $\mathcal{R}^{k}$ is the radiation resistance, and $\mathcal{X}^{k}$ is the modal reactance [23]. The current density modes are, in general, not orthogonal with the respect to the standard inner product on $\mathcal{B}$, defined in Eq. (29), but they are bi-orthogonal [24, 25], namely
@@
We now assume that the scattering response of each sphere is dominated by the electric dipole modes, each characterized by the impedance $\mathcal{Z}^{h}=\mathcal{Z}^{\mathrm{ed}}$, for $h=1,2,3$, where $\mathcal{Z}^{\mathrm{ed}}=\mathcal{R}^{\mathrm{ed}}+j\mathcal{X}^{\mathrm{ed}}$ is the electric dipole radiation impedance. The electric dipole radiation resistance ($\mathcal{R}^{\mathrm{ed}}$) and reactance ($\mathcal{X}^{\mathrm{ed}}$) are shown as a function of $k_{0}a$ in Fig. 2(a)-(b), respectively. This impedance is always ohmic–capacitive $(\mathcal{X}^{\mathrm{ed}}<0)$, and it can be expanded as in Eq. (46).
Figure 2: (a) Normalized electric dipole resistance $\mathcal{R}_{\mathrm{ed}}/\zeta_{0}$ (blue curve) and (b) reactance $\mathcal{X}_{\mathrm{ed}}/\zeta_{0}$ (red curve) as a function of $k_{0}a$. The RLC approximations (dashed line) are provided for reference. (c),(e) Normalized mutual resistance $R_{M}^{h}/(\zeta_{0}\rho^{2})$ and (d),(f) reactance $X_{M}^{h}/(\zeta_{0}\rho^{2})$ of the longitudinal (c)-(d) and transverse (e)-(f) coupling as a function of the electric distance $k_{0}d$. The first zeros of the mutual resistance are marked with $\times$.
@@
and $\rho=a/d$. The expressions of the mutual resistance $R_{M}^{h}=\Re\{Z_{M}^{h}\}$ and reactance $X_{M}^{h}=\Im\{Z_{M}^{h}\}$ are given in Eq. (48) and are plotted in Fig. 2(c)-(f) as functions of the electric interparticle distance $k_{0}d$, for both longitudinal and transverse polarizations. The oscillations describe in-phase and counter-phase radiative coupling.
@@
Notably, as shown in Appendix D, Eqs. (46a) and (49a), the leading-order terms in the expansions of $\mathcal{R}^{\mathrm{ed}}$ and $R_{M}^{h}$ are identical and approach zero in the quasistatic limit, defined by $k_{0}a\to 0$ and $k_{0}d\to 0$. Consequently, in the quasistatic limit, the circuit in Fig. 1(b) reduces to the simplified form in Fig. 1(c). In this limit, the only dissipative contribution arises from the two material resistances.
@@ III Resonances and Exceptional Points
The dynamical behavior of the circuit shown in Fig. 1(b) can be described by a linear time-invariant state-space model with state matrix $A$ [cf. Eq. (51)]. In the complex $\omega$-plane, the eigenfrequencies are related to the eigenvalues $\{\lambda_{i}\}$ of $A$ as $\omega_{i}=-j\lambda_{i}$.
The input impedance is defined as the impedance observed at one port when the remaining port is terminated by a short circuit, [see Fig. 1(d)]. As shown in the Appendix E, the zeros of the input impedance are eigenfrequencies of the system. The corresponding characteristic equation is then:
@@
However, in the full-wave regime, $\mathscr{P}\mathscr{T}$-symmetry of the circuit in Fig. 1(b) is generally broken because $R_{M}\neq 0$ almost-everywhere. Thus, in this regime, condition (17) no longer guarantees the validity of Eq. (16). Neverthless, it is easy to show that, if the following relation holds,
$Z_{\Sigma}(\omega;\boldsymbol{\vartheta})=\left[R_{M}+2jX_{M}\right]\frac{R_{M}}{Z_{22}},$ (18)
then, the characteristic equation (14) can be recast as
$\left|Z_{22}(\omega;\boldsymbol{\vartheta})\right|^{2}-\left(X_{M}(\omega;\boldsymbol{\vartheta})\right)^{2}=0,$ (19)
@@
Equation (20) represents one of the central results of this work: for a given geometry and material impedance $Z^{\mathrm{m}}_{2}$, it provides an explicit synthesis condition for $Z^{\mathrm{m}}_{1}$ that enforces operation at an exceptional point. Combined with Eq. (2), this condition can be directly recast in terms of the susceptibilities $\chi_{1}$ and $\chi_{2}$ of the two spheres.
As a special case, at the discrete values of $k_{0}d$ where the mutual resistance vanished (marked by $\times$ in Fig. 2(c),(e)), Eq. (18) gives $Z_{\Sigma}=0$. Consequently, (20) gives $R^{\mathrm{m}}_{1}=-(R^{\mathrm{m}}_{2}+2\mathcal{R}^{\mathrm{ed}})$ and $X^{\mathrm{m}}_{1}=X^{\mathrm{m}}_{2}$.
@@ IV Results
So far we have made no assumption about the dispersion relation of the material of the two spheres. In this section, we consider two spheres of radius $a$, whose susceptibility follows a Drude model, given by [29]:
@@
where $\omega_{\mathrm{p}\mu}\in\mathbb{R}$ is the plasma frequency, and $\gamma_{\mu}\in\mathbb{R}$ is the damping (or gain) rate associated with the free electron response in the $\mu$-th particle. For the Drude susceptibility (21), the material impedance corresponds to a simple RL series circuit; that is, $Z^{\mathrm{m}}_{\mu}(\omega)=R^{\mathrm{m}}_{\mu}+j\omega L^{\mathrm{m}}_{\mu}$, with $L^{\mathrm{m}}_{\mu}=(\omega_{\mathrm{p}\mu}^{2}\varepsilon_{0}a)^{-1}$ and $R^{\mathrm{m}}_{\mu}=\gamma_{\mu}L^{\mathrm{m}}_{\mu}$. Therefore, it may resonate with the capacitive radiation resistances of the dipole modes of the two spheres. For Drude susceptibility, the parameters are $\boldsymbol{\vartheta}=(\gamma_{1},\gamma_{2},k_{\mathrm{p}1}a,k_{\mathrm{p}2}a,d/a)$ where $k_{\mathrm{p}\mu}=\omega_{\mathrm{p}\mu}\sqrt{\varepsilon_{0}\mu_{0}}$.

### v2

@@ Abstract
We investigate exceptional points of degeneracy (EPDs) in electromagnetic scattering of a sphere dimer from the electroquasistatic limit to the fully retarded regime. In the quasistatic limit, we prove that $\mathscr{P}\mathscr{T}$-symmetric configurations, realized by spheres with complex-conjugate susceptibilities, host EPDs. Beyond this limit, retardation breaks $\mathscr{PT}$-symmetry; nevertheless, by jointly tuning the material dispersion of the two spheres, we derive analytic synthesis conditions for realizing EPDs at real frequencies. Near an EPD, we show that single-parameter perturbations yield the characteristic square-root splitting of the eigenfrequencies, and we quantify its impact on scattering, extinction, and absorption, clarifying sensing implications.
@@ I Introduction
Several studies [6, 7, 8] have employed coupled-mode theory (CMT) [13] to investigate EPDs in parity-time ($\mathscr{P}\mathscr{T}$) symmetric systems. While CMT is attractive because it is conceptually simple, offers clear physical intuition, and requires minimal computational effort, it does not precisely capture the underlying physics. In particular, coupled electromagnetic resonators are intrinsically non-$\mathscr{P}\mathscr{T}$ symmetric due to retardation in the radiative coupling, which leads to a complex and frequency-dependent coupling that cannot be easily included in conventional CMT. Moreover, a recent study [14], employing the coupled-dipole approximation (CDA) [15, 16, 17, 18, 19, 20, 21, 22, 23], has shown that these retardation effects can themselves induce EPDs even in systems with no gain and loss.
In this work, we investigate EPDs in electromagnetic scattering by a dimer of homogeneous, linear, time-dispersive spheres with arbitrary material dispersion. By using a volume-integral formulation of Maxwell’s equations, we recast the problem as an equivalent two-port circuit under the same assumptions as the coupled-dipole approximation [23, 22, 18]. Relative to the standard CDA, our formulation provides two main advantages: it disentangles the roles of excitation, geometry, material properties, radiation losses and scattering reactive fields, and it naturally leads to a state-space formulation. These features are essential for simplifying the analytical derivation and physical interpretation of EPD conditions. We prove that, in the electroquasistatic regime, a $\mathscr{P}\mathscr{T}$-symmetric sphere dimer with complex-conjugate susceptibilities supports EPDs. Beyond this limit, retardation breaks $\mathscr{P}\mathscr{T}$-symmetry; nevertheless, real-frequency EPDs can still emerge when the dispersions of the two materials are jointly tuned, and we derive general conditions for their occurrence. These predictions are validated against rigorous multiparticle Mie theory calculations [24]. Finally, we quantify the square-root spectral splitting to single-parameter perturbations and its impact on scattering, extinction, and absorption, highlighting the potential of this platform for highly sensitive sensing applications.
@@ II The Model
Figure 1: (a) Dimer composed of two spheres of radius $a$ and inter-center distance $d$. (b) Two-port equivalent circuit of two interacting particles in the coupled dipole approximation and in the full-wave regime. (c) Two-port equivalent circuit of two interacting particles in the coupled dipole approximation and in the static limit. (d) Configuration for characterizing the resonant mode of the dimer (i.e., $Z_{\mathrm{in}}=0$). In panels (b),(c),(d) $h\in\{x,y,z\}$ denotes the polarization of the dipolar mode.
Consider a dimer composed of two spherical particles, labeled by $\mu=1,2$, each made of a linear, homogeneous, isotropic, and time-dispersive material, characterized by an electric susceptibility $\chi_{\mu}$, embedded in vacuum. Each sphere has radius $a$ and is centered at $\mathbf{r}_{\mu}$, with the inter-center distance denoted by $d$. A time-harmonic dependence $e^{j\omega t}$ is assumed. Let $\mathcal{B}_{\mu}$ be the bounded three-dimensional domain occupied by the $\mu$-th particle, and $\partial\mathcal{B}_{\mu}$ its boundary. Without loss of generality, we assume that the dimer is aligned along the $x$-axis [see Fig. 1(a)].
A standard way to formulate the full-wave electromagnetic scattering problem, following the volume-integral-equation framework (see, e.g., [25, 26]), is to introduce as unknown the polarization current density $\mathbf{J}$ in the region $\mathcal{B}=\mathcal{B}_{1}\cup\mathcal{B}_{2}$ induced by the incident field $\mathbf{E}^{\mathrm{i}}$, which satisfies the homogeneous Helmholtz equation in vacuum. The volume-integral formulation enables a disentanglement of material response, radiative and reactive scattered fields, and geometry [27], crucial for deriving EPD synthesis conditions.
In the following, we adopt the notation described in Appendix A, where for any vector field $\mathbf{F}=\mathbf{F}(\mathbf{r})$ we denote by $\mathbf{F}_{\mu}$ its restriction on $\mathcal{B}_{\mu}$. In the region $\mathcal{B}_{\mu}$, the field ${\bf J}_{\mu}$ is related to the total electric field ${\mathbf{E}}_{\mu}$, which is the sum of the scattered field ${\mathbf{E}}_{\mu}^{\mathrm{s}}$ and the incident field ${\mathbf{E}}_{\mu}^{\mathrm{i}}$, i.e. $\mathbf{E}_{\mu}=\mathbf{E}_{\mu}^{\mathrm{i}}+\mathbf{E}_{\mu}^{\mathrm{s}}$, by the constitutive relation
@@
where $Z_{\mu}^{\mathrm{m}}$ is the material impedance of the $\mu$-th particle [28, 29], defined as
@@
where $\mathbf{E}^{\mathrm{s}}_{\mu,\nu}$ denotes the contribution to the scattered field on $\mathcal{B}_{\mu}$ arising from the current density $\mathbf{J}_{\nu}$ on the $\nu$-th particle alone. The scattered field $\mathbf{E}^{\mathrm{s}}_{\mu,\nu}$ is related to its source $\mathbf{J}_{\nu}$ through the volume integral operator $\mathscr{L}_{\mu}$ as
@@
where $I_{\mu}^{k}=I_{\mu}^{k}(\omega)$ are the weights. The polarization current density modes $\mathbf{j}^{k}_{\mu}$ are the solutions of the following eigenvalue problem in each individual bounded domain $\mathcal{B}_{\mu}$:
$\mathscr{L}_{\mu}\left\{\mathbf{j}^{k}_{\mu}\right\}=\mathcal{Z}^{k}\,\mathbf{j}^{k}_{\mu},\qquad\mathbf{r}\in\mathcal{B}_{\mu}$ (6)
where $\mathcal{Z}^{k}=\mathcal{R}^{k}+j\mathcal{X}^{k}$ is the mode impedance of the current eigenmodes $\mathbf{j}_{\mu}^{k}$, $\mathcal{R}^{k}$ is the radiation resistance, and $\mathcal{X}^{k}$ is the modal reactance [28, 29]. The current density modes have support within the particle volume and are, in general, not orthogonal with respect to the standard inner product on $\mathcal{B}$, defined in Eq. (29). However, they satisfy the bi-orthogonality relation [30, 31]:
@@
We now assume that the scattering response of each sphere is dominated by the electric-dipole modes, each characterized by the impedance $\mathcal{Z}^{h}=\mathcal{Z}^{\mathrm{ed}}$, for $h=1,2,3$, where $\mathcal{Z}^{\mathrm{ed}}=\mathcal{R}^{\mathrm{ed}}+j\mathcal{X}^{\mathrm{ed}}$ is the electric-dipole impedance. The electric-dipole radiation resistance ($\mathcal{R}^{\mathrm{ed}}$) and reactance ($\mathcal{X}^{\mathrm{ed}}$) are shown as a function of $k_{0}a$ in Fig. 2(a)-(b), respectively. This impedance is always ohmic–capacitive $(\mathcal{X}^{\mathrm{ed}}<0)$, and it can be expanded as in Eq. (46).
Figure 2: (a) Normalized electric-dipole resistance $\mathcal{R}^{\mathrm{ed}}/\zeta_{0}$ (blue curve) and (b) reactance $\mathcal{X}^{\mathrm{ed}}/\zeta_{0}$ (red curve) as a function of $k_{0}a$. The RLC approximations in Eq. (46) (dashed lines) are provided for reference. (c),(e) Normalized mutual resistance $R_{M}^{h}/(\zeta_{0}\rho^{2})$ and (d),(f) reactance $X_{M}^{h}/(\zeta_{0}\rho^{2})$ of the longitudinal (c)-(d) and transverse (e)-(f) coupling as a function of the electric distance $k_{0}d$. The first zeros of the mutual resistance are marked with $\times$.
@@
and $\rho=a/d$. The expressions of the mutual resistance $R_{M}^{h}=\operatorname{Re}\{Z_{M}^{h}\}$ and reactance $X_{M}^{h}=\operatorname{Im}\{Z_{M}^{h}\}$ are given in Eq. (48) and are plotted in Fig. 2(c)-(f) as functions of the electric interparticle distance $k_{0}d$, for both longitudinal $(h=x)$ and transverse polarizations $(h=y,z)$. The oscillations describe in-phase and counter-phase radiative coupling.
@@
Notably, as shown in Appendix D, Eqs. (46a) and (49a), the leading-order terms in the expansions of $\mathcal{R}^{\mathrm{ed}}$ and $R_{M}^{h}$ are identical and approach zero in the quasistatic limit, defined by $k_{0}a\to 0$ and $k_{0}d\to 0$. In this limit, the electric-dipole impedance and the mutual impedance reduce to their leading capacitive terms, namely $\mathcal{Z}^{\mathrm{ed}}\sim 1/(j\omega C^{\mathrm{ed}})$, and $Z_{M}^{h}\sim 1/(j\omega C_{M}^{h})$. Consequently, in the quasistatic limit, the circuit in Fig. 1(b) reduces to the simplified form in Fig. 1(c). In this limit, the only dissipative contribution arises from the two material resistances.
@@ III Resonances and Exceptional Points
The dynamical behavior of the circuit shown in Fig. 1(b) can be described by a linear time-invariant state-space model with state matrix $\boldsymbol{A}$ [cf. Eq. (51)]. In the complex $\omega$-plane, the eigenfrequencies are related to the eigenvalues $\{\lambda_{i}\}$ of $\boldsymbol{A}$ as $\omega_{i}=-j\lambda_{i}$.
The input impedance is defined as the impedance observed at one port when the remaining port is terminated by a short circuit [see Fig. 1(d)]. As shown in Appendix E, the zeros of the input impedance are eigenfrequencies of the system. The corresponding characteristic equation is then:
@@
However, in the full-wave regime, $\mathscr{P}\mathscr{T}$-symmetry of the circuit in Fig. 1(b) is generally broken because $R_{M}\neq 0$ almost everywhere. Thus, in this regime, condition (17) no longer guarantees the validity of Eq. (16). Nevertheless, it is easy to show that, if the following relation holds,
$Z_{\Sigma}(\omega;\boldsymbol{\vartheta})=\left(R_{M}+2jX_{M}\right)\frac{R_{M}}{Z_{22}},$ (18)
then the characteristic equation (14) can be recast as
$\left|Z_{22}(\omega;\boldsymbol{\vartheta})\right|^{2}-X_{M}^{2}(\omega;\boldsymbol{\vartheta})=0,$ (19)
@@
Equation (20) represents one of the central results of this work: for a given geometry and material impedance $Z^{\mathrm{m}}_{2}$, it provides an explicit synthesis condition for $Z^{\mathrm{m}}_{1}$ which can be directly recast in terms of the susceptibilities $\chi_{1}$ and $\chi_{2}$ through Eq. (2).
As a special case, at the discrete values of $k_{0}d$ where the mutual resistance vanishes (marked by $\times$ in Fig. 2(c),(e)), Eq. (18) gives $Z_{\Sigma}=0$. Consequently, (20) gives $R^{\mathrm{m}}_{1}=-(R^{\mathrm{m}}_{2}+2\mathcal{R}^{\mathrm{ed}})$ and $X^{\mathrm{m}}_{1}=X^{\mathrm{m}}_{2}$.
@@ IV Results
So far we have made no assumption about the material dispersion relation of the two spheres. In this section, we consider two spheres of radius $a$, whose susceptibility follows a Drude model, given by [35]:
@@
where $\omega_{\mathrm{p}\mu}\in\mathbb{R}$ is the plasma frequency, and $\gamma_{\mu}\in\mathbb{R}$ is the damping (or gain) rate associated with the free electron response in the $\mu$-th particle. For the Drude susceptibility (21), the material impedance corresponds to a simple RL series circuit; that is, $Z^{\mathrm{m}}_{\mu}(\omega)=R^{\mathrm{m}}_{\mu}+j\omega L^{\mathrm{m}}_{\mu}$, with $L^{\mathrm{m}}_{\mu}=(\omega_{\mathrm{p}\mu}^{2}\varepsilon_{0}a)^{-1}$ and $R^{\mathrm{m}}_{\mu}=\gamma_{\mu}L^{\mathrm{m}}_{\mu}$. Therefore, it can resonate with the capacitive impedances of the dipole modes of the two spheres. For the Drude susceptibility, the parameters are $\boldsymbol{\vartheta}=(\gamma_{1},\gamma_{2},k_{\mathrm{p}1}a,k_{\mathrm{p}2}a,d/a)$ where $k_{\mathrm{p}\mu}=\omega_{\mathrm{p}\mu}\sqrt{\varepsilon_{0}\mu_{0}}$.